OpenStudy (anonymous):

vertices - 14 edges - 36 faces? ????? medal :) thanks

OpenStudy (ivettef365):

Use formula F + V - E = 2 F +14 - 36 = 2 F 22 = 2 F = 20

OpenStudy (ivettef365):

F - 22 = 2 F = 20